OpenStudy (anonymous):

Derivative Applications in Motion and Economics Velocity and Acceleration A ball is dropped from at height of 100 ft. One second later another ball is dropped from a height of 75 ft. which one hits the ground first?

OpenStudy (anonymous):

alright this is the way I did it. I graphed on the ti these two equations (y=-16t^2+100, and y=-16t^2+75) the one from 75 ft crosses the x axis closer to the point of origin so i'm going with that one hitting the ground faster

OpenStudy (anonymous):

or you could solve for t in \[-16t^2+100=0\] and \[-16t^2+75=0\] then add 1 to the second answer and compare it to the first